OpenStudy (henryarias14):

find derivative of x9^x

OpenStudy (henryarias14):

please explain steps too

OpenStudy (tkhunny):

Go! Looks like you have a product rule on your hands. Let's see your best work.

OpenStudy (henryarias14):

it says i should rewrite using logarithmic properties before differentiating

OpenStudy (henryarias14):

oh one sec let me give it a try

OpenStudy (henryarias14):

product rule: 9^x + xln9(9^x)

OpenStudy (henryarias14):

right?

OpenStudy (tkhunny):

I don't know why you would do that for this one, but okay. y = x9^x Then log y = log(x9^x) = log(x) + log(9^x) = log(x) + xlog(9) Now, you have an implicit definition, requiring the Chain Rule on the left. The right is pretty straight-forward.

OpenStudy (henryarias14):

1/yln(10)?

OpenStudy (henryarias14):

book has 9^x(xln9+1)

OpenStudy (peachpi):

you got the book answer with the product rule. To pick up from tk: \[y = x9^x\] \[\ln y=\ln(x9^x)\] \[\ln y = \ln x + \ln 9^x\] \[\ln y = \ln x + x \ln 9\] \[\frac{ 1 }{ y }\frac{ dy }{ dx }=\frac{1}{x}+ln 9\] \[\frac{ dy }{ dx }=y \left( \frac{ 1 }{ x }+\ln 9 \right)\] \[\frac{ dy }{ dx }=x9^x \left( \frac{ 1 }{ x }+\ln 9 \right)\] \[\frac{ dy }{ dx }=9^x \left( 1+x \ln 9 \right)\] I really don't see why they'd have you do it this way, unless it was just for practice.

OpenStudy (tkhunny):

That was my impression. Practice introducing the logarithm. Note to OP, You WILL need the logarithm for derivatives like \(\dfrac{d}{dx}x^{x}\) or limits like \(\underset{{x\to \infty}}{\lim}\left(1 + \dfrac{a}{x}\right)^{x}\). With the variable in both Base and Exponent, it takes just a bit more effort.